Write in center-radius form. Give the center and radius. Hint: Use completing the square. x^2 y^2-10x-6y 25=0
Answer by lwsshak3(11628)

Write in center-radius form. Give the center and radius. Hint: Use completing the square.
x^2+ y^2-10x-6y+ 25=0
**
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k) being the (x,y) coordinates of the center and r=radius
..
complete the squares
(x^2-10x+25)+(y^2-6y+9)=-25+25+9=9
equation:
(x-5)^2+(y-3)^2=9
Center: (5/2)
Radius=√9=3